A hat contains 20 names , 12 of which are female . If five names are drawn from the hat (without replacement) , what is the probability that there are at least two female names drawn ?
1
Expert's answer
2020-03-16T14:50:10-0400

"P(x\geq 2)" is the probability that there are at least two female names drawn

"P(x\geq 2)=1-P(x<2)=1-P(x=1)-P(x=0)"


"_{20}C_5" is the number of ways we can choose 5 names from 20.


"P(x=1)" is the probability that there is one female name drawn, so there are "_{12}C_1" ways to choose 1 female name from 12, and there are "_8C_4" ways to choose other 4 male names from 8.

"P(x=1)=\frac{_{12}C_1 \times _8 C_4}{_{20} C_5}=\frac{12\times 70}{15504}=\frac{35}{646}"


"P(x=0)" is the probability that there is no female names drawn, so all 5 names are male, we can choose them in "_8C_5" ways.

"P(x=0)= \frac{_8 C_5}{_{20} C_5}=\frac{56}{15504}=\frac{7}{1938}"


"P(x\geq 2)=1-\frac{35}{646}-\frac{7}{1938}=\frac{913}{969}=0.942"


Answer: 0.942